Project journal

Title Description Author Post date
Weekly Update 0/5/08/16

All,

Please find the weekly status update.

RAG Status: Amber

Activity since last update:

  • Process walkthrough complete.
  • SIP trunks (telephony channels) secured.
  • Agent licences secured.
  • Headsets ordered (after recent decision to use these instead of handsets - costs approved SRA).
  • Call flow structure complete.
  • Phone numbers allocated.
  • Deployment of kit to call centre initiated.

Planned

Iain Moore-McLaughlin 11-Aug-2016
Weekly Update 22/07/2016

RAG STATUS - Amber

Activity since last update:

•             Risks discussed and assigned to relevant personnel.

•             Contact made with relevant desktop support personnel and discussions planned for on site support.

•             Telephony solution in progress and looking good for completion.

Planned

•             Approval of telephony proposal detailing what can/cannot be done given existing technology and current systems.

•             Commencement of telephony solution configuration.

Iain Moore-McLaughlin 27-Jul-2016
Weekly Update 29/07/2016

All,

 

Please find the weekly status update.

RAG Status: Amber

Activity since last update:

  • Telephony Solution developed and submitted for approval (approval expected ASAP).
  • Tier 1 staff recruitment in progress.
  • Go day Telephony/Desktop support secured.

Planned

  • Process walkthrough meeting to ensure operations will be streamlined on day.
  • Deploy phones to call centre site.

Key Risks

Iain Moore-McLaughlin 29-Jul-2016
Weekly Update 15/07/2016

RAG STATUS - Amber

Activity since last update:

•             Further planning meeting went ahead including college admin staff now all returned from leave.

•             Site with adequate space, kit and network points secured to house all involved staff if required.

•             Plan created but still requires some fine tuning.

Planned

•             Creation of telephony proposal detailing what can/cannot be done given existing technology and current systems.

Iain Moore-McLaughlin 15-Jul-2016
Weekly Update 07/07/2016

RAG STATUS - Green

Activity So Far

* Kick off meeting went ahead as planned on Tuesday.

* Activity commenced to nail down where College Admissions are with activity required of them.

* High level view of outputs established (e.g. potential kit requirements, call scripts, support reqs. processes etc.).

* Activity commenced on a brief technical proposal taking Ian Sutherlands call flow diagram into consideration.

Iain Moore-McLaughlin 15-Jul-2016